OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. 48, 8, four divided by three, two divided by nine, ... @agent0smith

OpenStudy (agent0smith):

First: Find the common ratio, by dividing a term by the term before it (use the 2nd and 1st term)

OpenStudy (anonymous):

0.16666

OpenStudy (agent0smith):

So it converges since the common ratio is smaller than 1, now find the sum \[\Large \frac{ a_1 }{ 1-r }\]

OpenStudy (anonymous):

Wait confused haha

OpenStudy (agent0smith):

By...?

OpenStudy (anonymous):

IS A1 0.1666

OpenStudy (agent0smith):

a1 is always first term. r is common ratio.

OpenStudy (anonymous):

47.84

OpenStudy (agent0smith):

Can't be right... a1 is 48 and r is 1/6, the sum will be greater than 48.

OpenStudy (anonymous):

I keep getting 47 >.<

OpenStudy (agent0smith):

48/(1 - 1/6) =

OpenStudy (anonymous):

57.6

OpenStudy (agent0smith):

Yep :)

OpenStudy (anonymous):

But okay what do i do now? Determine whether the sequence converges or diverges. If it converges, give the limit.

OpenStudy (agent0smith):

We've done everything already.

OpenStudy (anonymous):

but these are the aswer choiced : Converges; two hundred and eighty eight divided by five Converges; 0 Diverges Converges; -12432

OpenStudy (agent0smith):

check this on a calculator: two hundred and eighty eight divided by five

OpenStudy (anonymous):

56.6

OpenStudy (agent0smith):

288/5 = 57.6 but yep :)

OpenStudy (anonymous):

Find an equation for the nth term of the sequence. -3, -12, -48, -192, ...

OpenStudy (agent0smith):

Find the common ratio same way as before. First term is easy. Then plug them in \[\huge a_n = a_1 r^{n-1}\]

OpenStudy (anonymous):

4 is the common ratio

OpenStudy (agent0smith):

Yes and a1 is...?

OpenStudy (anonymous):

3?

OpenStudy (agent0smith):

a1 is just the first term

OpenStudy (anonymous):

so -3

OpenStudy (anonymous):

an = -3 • 4n - 1

OpenStudy (agent0smith):

yep, but make sure the (n-1) is the exponent

OpenStudy (anonymous):

Find an equation for the nth term of a geometric sequence where the second and fifth terms are -21 and 567, respectively.

OpenStudy (agent0smith):

Make a new question for this one, it needs more work